OpenStudy (anonymous):

HELP!!!!! 2z + 5 + 3z a. 10z c. 5z + 5 b. 7z + 3 d. 8z + 2 i think its c. but then again its prolly a. i need help!!

OpenStudy (ivettef365):

combine like terms

hartnn (hartnn):

2z+3z = (2+3)z=5z and you cannot combine 5z and 5 because they are not like terms